OpenStudy (anonymous):

Ruben wants to decorate the edge of his circular flower garden with some fancy edge trim. The diameter of the garden is 15 feet. If the edge trp costs 3 1/2 dollars per linear foot, how much would it cost to edge trim the garden? (use pie and round to the nearest cent)

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

hint: the length of the edge of the flower garden is: \[\large L = 2 \times \pi \times R\] where: R is the radius, namely: \[\large R = \frac{D}{2}\] and D is the diameter

OpenStudy (gorv):

edge means circumference of the circle here

OpenStudy (anonymous):

L=2*3.14*15?

OpenStudy (michele_laino):

no, L=2*3.14*7.5 since the radius is: R = 15/2=7.5 feet

OpenStudy (anonymous):

oh gotchya! k now what next step? I got 47.1 for the above part

OpenStudy (michele_laino):

ok!

OpenStudy (michele_laino):

now, we have to multiply that length by 3 1/2

OpenStudy (anonymous):

okay and that gave me 164.85

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

oh yay!!!

OpenStudy (michele_laino):

:)